# Harpa: High-Rate Phase Association with Travel Time Neural Fields

### v2

@@
Harpa: High-Rate Phase Association with Travel Time Neural Fields
Cheng Shi ${}^{1}$ Maarten V. de Hoop ${}^{2}$ Ivan Dokmanić${}^{1,3\ast}$
${}^{1}$epartement of Mathematics and Computer Science, University of Basel,
${}^{2}$imons Chair in Computational and Applied Mathematics and Earth Science, Rice University
${}^{3}$epartment of ECE, University of Illinois at Urbana–Champaign
${}^{\ast}$o whom correspondence should be addressed; E-mail: ivan.dokmanic@unibas.ch
@@
Teaser
Physics-informed generative ML enables very high rate seismic phase association despite large wavespeed uncertainty.
Introduction
Seismic signals are a gateway to understanding the structure and dynamics of the Earth. Recent breakthroughs in deep-learning-based seismic processing enable reliable detection of very small, high-rate seismic events (?, ?, ?) with potential to reveal unprecedented details about the variability of elastic material properties in the Earth’s crust and upper mantle. Progress is, however, stymied by the limitations of existing phase association methodologies designed for larger events which are less frequent, following the Gutenberg–Richter law. High arrival rates such as those associated with seismic clouds (?) pose unique challenges.
@@ Introduction
To address this gap we propose Harpa—a method that can solve the high-rate association problem over an unknown wave speed in settings considerably more complex than those addressed by existing methods. Harpa first solves a seemingly harder problem: estimating locations and occurrence times of earthquakes jointly with an (approximate) wave speed model.
@@
Figure 1: Seismic events on different time scales and magnitudes: Panel a: seismic data from Ridgecrest CA, recorded on July 6, 2019, between 04:10 and 04:30. It has both large, isolated events and small, clustered ones. Triangle sizes represent earthquake magnitudes; color represents arrival times. The density distribution of magnitudes illustrated in the inset plot signifies the Gutenberg-Richter law where the slope represents the $b$ value. Panel b: low-rate large magnitude events which are easy to associate. Each trace on the plot corresponds to a recording from a station, which is positioned according to its respective latitude value on the y-axis. The diameters of the disks are indicative of the peak ground velocity (PGV) (?). Color represents the association, that is, waveforms from a common seismic source. Gray disks in c, d indicate the unidentified arrivals in (?). Large magnitude arrivals in panel b are straightforwardedly linked. Panel c illustrates a time interval, at finer scale, that is more challenging for association but nonetheless successful as the arrival time fluctuations due to the henetrogeneity in wavespeed are still small compared to the intermittancy (rate) of events. At yet a finer time scale and shorter interval in panel d identification and association becomes a challenge and has, by traditional techniques, been unresolved. This intermittancy or rates originate from yet smaller earthquakes in accordance with the Gutenberg-Richter law. On the one hand, the detection of ever smaller events is challenging due to “background” noise; on the other hand, the implied increasing rate poses the challenge of association due to wavespeed fluctuations. It is the latter that we address and illustrate in panels e, f, and g. Here, we demonstrate this using a synthetic wave speed model, the SEG/EAGE 3-D overthrust, with congruent spatial dimensions. Blue and orange triangles indicate two earthquakes occurring 20 seconds apart, with easily associated arrivals (panel g). The remaining four smaller triangles indicate events which all occur within 2 seconds, presenting a significantly harder association challenge where the ordinal number of arrival carries no significance (panel g). It is worthwhile noting that under stronger wave speed fluctuations the association challenge addressed here appears at larger magnitude earthquakes.
Only then does it assign earthquake phases, but with the estimated earthquake and wave speed parameters the problem reduces to a simpler linear assignment problem. This simplification is afforded by leveraging the physics of wave propagation and advances in generative modeling and optimal transport.
@@
There are two major challenges in implementing this program. The first is to enable numerically efficient differentiable optimization with respect to source locations and wave speed models which respects propagation physics. This is difficult with traditional mesh-based eikonal solvers; scaling to 3D would result in prohibitive computational complexity (?). The second challenge lies in finding the globally optimal solution of the associated non-convex optimization problem. While the problem of wave speed reconstruction from unknown sources admits a unique solution under certain conditions (?), identifying this solution by optimizing a non-convex objective is challenging due to the presence of poor local minima.
To resolve the first challenge, we first train a deep generative model to embed a family of complex wave speeds in a low-dimensional latent space and optimize the latent code of the wave speed and the spatio-temporal earthquake parameters simultaneously. But instead of directly addressing the formidable challenge of optimizing these parameters through a numerical solver we train a deep neural field (?, ?, ?) to fit the travel times between arbitrary interior points and a fixed set of receivers. An autoencoder structures the latent space and provides a pairing between latent codes and wave speeds which is then used by the travel time neural field. As we show in Experiments, fitting the best wave speed from the range of a sufficiently diverse generative model makes the association highly robust, even when the test wave speed is far out of the distribution used to train the model.
To resolve the second challenge, we leverage a combination of stochastic gradient-based methods and sampling. Markov Chain Monte Carlo (MCMC) methods can in principle deal with the non-convexity but they are ineffective in our setting because the search space is too large. We employ Stochastic Gradient Langevin Dynamics (SGLD) (?) to ensure an adequate exploration of the loss landscape by the gradient optimizer, ultimately leading to the identification of a (near-)global optimum.
@@
Background and related work
If earthquake source locations, occurrence times, and wave speeds are all known, the phase association problem can be formulated as a linear sum assignment problem (LSAP), which is effectively solved by the Hungarian algorithm (?). If only the occurrence times are unknown, the problem is related to a quadratic assignment problem (QAP) (?, ?) which is NP-hard. While QAPs can be solved in certain practical cases using heuristics and meta-heuristics (?, ?), there has been growing interest in leveraging machine learning to solve similar combinatorial optimization problems (?, ?, ?). The situation, however, is often even more complex: neither the wave speed nor the spatiotemporal earthquake source locations are known, although sometimes a rather coarse wave speed model is available. These unknown degrees of freedom result in a mixed-integer programming problem which is harder than LSAP or QAP and challenging to solve even with a small number of earthquakes and stations (?, ?, ?).
Although the joint localization–wave speed recovery problem has a unique solution under certain conditions (?), there are no reliable algorithms that solve it provably and efficiently. Notwithstanding, there exist algorithmic attempts to perform joint full waveform (or sometimes travel time) inversion and wave speed model refinement but with temporally well-separated sources, thus implicitly avoiding the association problem (while still facing the ill-posedness) (?, ?, ?). In the context of micro-seismicity, simple parametric models (?) and Bayesian priors (?) were used for regularization. We mention that our strategy to convert combinatorial data into measures in order to avoid combinatorial optimization is reminiscent of earlier work on unassigned Euclidean distance data (?, ?) with applications in genomics and room acoustics.
Over the last decade there has been an increasing interest in the application of machine learning in seismology, for example for phase picking (?, ?, ?, ?). A multitude of studies addressing the phase association problem have employed either probabilistic (?, ?, ?) or graph models (?, ?). Some adopt a static or 1-D wave speed model, while others incorporate wave speed information indirectly through supervised learning. Since these approaches do not explicitly leverage the wave propagation physics they demand large quantities of data for training. All phenomena in the association problem are governed by the wave equation which is known, but is typically only leveraged for downstream analysis once the phases have been associated. We show that rethinking the problem in measure-theoretic terms allows one to leverage the wave propagation model from the outset.
Results
Problem formulation
@@ Results
A mixed-integer optimization problem
@@
Generative modeling of low-dimensional wave speed families
To build Harpa we assume that the wave speed models and consequently travel times exhibit a latent low-dimensional structure. We thus model the wave speeds by an autoencoder-based generative model and traveltimes by a matching continuous coordinate-based neural field (?, ?, ?). The assumption of latent low-dimensionality is consistent with geodynamical models which generate the distribution of the elastic material properties in the interior; this has been explored for Earth’s mantle before (?, ?). Geological random fields have been used to in conjuction with accurate earthquake simulations to study the 2019 Le Teil earthquake in a region where little geophysical data is available (?). In fact, the wave speeds generated by our latent model ($\mathbb{R}^{L}\ni z\mapsto c(z)$) are considerably more complex than the fixed models used for association in the recent literature which only vary along the depth coordinate (?, ?, ?). We thus define the associated travel time function as $T(s,z;r)=T_{c}(s;r)$. This enables us to rewrite the optimization problem involving $c$ as one involving the latent code $z$,
@@
Association without association
@@
In the case where some phase picks are missing or spurious picks are present, we define a loss for matching unbalanced arrivals based on sparsity-constrained optimal transport; a detailed description is given in Materials and method and results in Fig. 9c. In the 1-dimensional discrete case this loss coincides with the Wasserstein loss of the unbalanced linear assignment.
@@
where *$\mathcal{E}^{*},z^{*}$ on the left hand side are the solution of (10) while $c^{*},\alpha^{*},(s^{*},\tau^{*})$ on the right hand side are the solution of (2). After obtaining *$\mathcal{E}^{*}=(s^{*},\tau^{*})$ from (10), we can find $\alpha^{*}$ by solving LSAP directly, or by using only $s^{*}$ without $\tau^{*}$ by solving the TDoA problem as in (4)).
We assume that earthquakes are well-approximated by point sources which are distributed volumetrically so that we obtain sufficiently rich data. These conditions are satisfied in reality with events occurring close to ensembles of faults (or fractures) at varying orientations, or in “clouds” (?, ?).
@@
Generative travel time neural fields
@@
The way we resolve this conundrum in Harpa is as follows: We first use a deep autoencoder to build a low-dimensional latent representation of a set of complex wave speeds; see Experiments and Fig. 3. Next, we use the fast marching method as implemented in scikit-fmm to compute the travel times between a fixed set of receivers and sources at each possible mesh node. The final and key step is then to fit a neural field (also known as an implicit neural representation) to the travel time data (?, ?). We use a neural network $T_{\theta}$ with weights $\theta$ and periodic activation functions (?) which ensures that travel times and their derivatives with respect to source locations and wave speed latent codes can be efficiently evaluated at continuous coordinates. We train $T_{\theta}$ so that $T_{\theta^{*}}(s,z;r)\approx T(s,z;r)$ fits the travel time at a discrete set of grid points in space,
@@
The Wasserstein loss in (10) is differentiable with respect to both $z$ and $\mathcal{E}$ but it is not convex in these parameters which means that local gradient-based optimization may converge to local minima. The total number of parameters, $(3+1)\times N+L$, is small compared with common deep neural networks so there is no reason to expect that these local minima are benign. Indeed, as shown in Materials and method, we observe empirically that stochastic gradient descent (SGD) often converges to poor local minima in this setting. Monte Carlo as an alternative would be prohibitively slow.
There exist modern iterative optimizers which combine the favorable aspects of sampling methods and SGD. A notable example is stochastic gradient Langevin dynamics (SGLD) (?), which can be viewed either as a mini-batch variant of Langevin dynamics or as a noisy SGD. The global convergence of Langevin dynamics can be leveraged to locate the globally-optimal earthquake and wave speed parameters which approximate their corresponding ground truth values, while benefiting from fast convergence rates (?, ?); here the (relatively) small number of parameters helps. A SGLD iterate is computed as
@@
Experiments
We now showcase the performance of Harpa on a suite of computer experiments; we work with both known and unknown wave speeds. In Association via joint recovery we first test the limits of Harpa on a challenging set of unknown wave speeds in terms of the numbers of sources and receivers and show successful association in a variety of settings which were heretofore inaccessible for association. In Distorted and out-of-distribution wave speeds we apply Harpa to realistic wave speed models derived from the SEG / EAGE 3-D overthrust model (?) which come from a different distribution than the one used to train the latent wave speed model and show that the proposed method is robust to such perturbations. Finally, in Spurious and missing picks we demonstrate robustness to spurious and missing phase picks.
Confusion factor and association evaluation
@@
calculates Kendall’s $\tau$ rank correlation coefficient (?). Succinctly, when the order of the arrivals at a station matches the order of earthquake occurrences, $\mathcal{T}$ yields a value of $1$ and $\mathsf{CF}=0$. This typically occurs during low-rate earthquakes, where the travel time is shorter than the time between the different events. Conversely, when all earthquakes occur almost simultaneously, the order of earthquake occurrence bears no correlation with the order of the phases/picks arriving at the stations. This reduces $\mathcal{T}$ closer to 00, and pushes $\mathsf{CF}$ closer to $1$, implying a higher level of confusion (the $\mathsf{CF}$ is always between 0 and 1). When studying regions such as Ridgecrest around the time of the 2019 earthquake, the magnitude cutoff is usually chosen so that the resulting CF is below 0.1 (?, ?, ?). Fig. 4 shows representative arrival configurations for different values of $\mathsf{CF}$; Fig. 9 shows results of an exhaustive test of the proposed method at different $\mathsf{CF}$.
The difficulty of association is not solely determined by $\mathsf{CF}$; it also depends on the number of stations $N$ and the number of sources $M$ being simultaneously associated. Many algorithms initially partition the data into shorter temporal windows, usually between 5 and 20 seconds in length, with approximately 5 to 20 events in each window (?, ?). In this proof of concept work we focus on a single window.
@@
Association via joint recovery

### v3

# High-Rate Phase Association with Travel Time Neural Fields

@@
Abstract
@@
1 Introduction
Seismic signals are a gateway to understanding the structure and dynamics of the Earth. Recent breakthroughs in deep-learning-based seismic processing enable reliable detection of very small, high-rate seismic events [1, 2, 3] with potential to reveal unprecedented details about the variability of elastic material properties in the Earth’s crust and upper mantle. Progress is, however, stymied by the limitations of existing phase association methodologies designed for larger events which are less frequent, following the Gutenberg–Richter law. High arrival rates such as those associated with seismic clouds [4] pose unique challenges.
@@ 1 Introduction
To address this gap we propose Harpa—a method that can solve the high-rate association problem over an unknown wave speed in settings considerably more complex than those addressed by existing methods. Harpa first solves a seemingly harder problem: estimating locations and occurrence times of earthquakes jointly with an (approximate) wave speed model. Only then does it assign earthquake phases, but with the estimated earthquake and wave speed parameters the problem reduces to a simpler linear assignment problem. This simplification is afforded by leveraging the physics of wave propagation and advances in generative modeling and optimal transport.
@@
Figure 1: Seismic events on different time scales and magnitudes: Panel a: seismic data from Ridgecrest CA, recorded on July 6, 2019, between 04:10 and 04:30. It has both large, isolated events and small, clustered ones. Triangle sizes represent earthquake magnitudes; color represents arrival times. The density distribution of magnitudes illustrated in the inset plot signifies the Gutenberg-Richter law where the slope represents the $b$ value. Panel b: low-rate large magnitude events which are easy to associate. Each trace on the plot corresponds to a recording from a station, which is positioned according to its respective latitude value on the y-axis. The diameters of the disks are indicative of the peak ground velocity (PGV) [3]. Color represents the association, that is, waveforms from a common seismic source. Gray disks in c, d indicate the unidentified arrivals in [5]. Large magnitude arrivals in panel b are straightforwardedly linked. Panel c illustrates a time interval, at finer scale, that is more challenging for association but nonetheless successful as the arrival time fluctuations due to the henetrogeneity in wavespeed are still small compared to the intermittancy (rate) of events. At yet a finer time scale and shorter interval in panel d identification and association becomes a challenge and has, by traditional techniques, been unresolved. This intermittancy or rates originate from yet smaller earthquakes in accordance with the Gutenberg-Richter law. On the one hand, the detection of ever smaller events is challenging due to “background” noise; on the other hand, the implied increasing rate poses the challenge of association due to wavespeed fluctuations. It is the latter that we address and illustrate in panels e, f, and g. Here, we demonstrate this using a synthetic wave speed model, the SEG/EAGE 3-D overthrust, with congruent spatial dimensions. Blue and orange triangles indicate two earthquakes occurring 20 seconds apart, with easily associated arrivals (panel g). The remaining four smaller triangles indicate events which all occur within 2 seconds, presenting a significantly harder association challenge where the ordinal number of arrival carries no significance (panel g). It is worthwhile noting that under stronger wave speed fluctuations the association challenge addressed here appears at larger magnitude earthquakes.
@@
There are two major challenges in implementing this program. The first is to enable numerically efficient differentiable optimization with respect to source locations and wave speed models which respects propagation physics. This is difficult with traditional mesh-based eikonal solvers; scaling to 3D would result in prohibitive computational complexity[6]. The second challenge lies in finding the globally optimal solution of the associated non-convex optimization problem. While the problem of wave speed reconstruction from unknown sources admits a unique solution under certain conditions [7], identifying this solution by optimizing a non-convex objective is challenging due to the presence of poor local minima.
To resolve the first challenge, we first train a deep generative model to embed a family of complex wave speeds in a low-dimensional latent space and optimize the latent code of the wave speed and the spatio-temporal earthquake parameters simultaneously. But instead of directly addressing the formidable challenge of optimizing these parameters through a numerical solver we train a deep neural field [8, 9, 10] to fit the travel times between arbitrary interior points and a fixed set of receivers. An autoencoder structures the latent space and provides a pairing between latent codes and wave speeds which is then used by the travel time neural field. As we show in Section 3, fitting the best wave speed from the range of a sufficiently diverse generative model makes the association highly robust, even when the test wave speed is far out of the distribution used to train the model.
To resolve the second challenge, we leverage a combination of stochastic gradient-based methods and sampling. Markov Chain Monte Carlo (MCMC) methods can in principle deal with the non-convexity but they are ineffective in our setting because the search space is too large. We employ Stochastic Gradient Langevin Dynamics (SGLD) [11] to ensure an adequate exploration of the loss landscape by the gradient optimizer, ultimately leading to the identification of a (near-)global optimum.
@@
1.1 Background and related work
If earthquake source locations, occurrence times, and wave speeds are all known, the phase association problem can be formulated as a linear sum assignment problem (LSAP), which is effectively solved by the Hungarian algorithm [12]. If only the occurrence times are unknown, the problem is related to a quadratic assignment problem (QAP) [13, 14] which is NP-hard. While QAPs can be solved in certain practical cases using heuristics and meta-heuristics [15, 16], there has been growing interest in leveraging machine learning to solve similar combinatorial optimization problems [17, 18, 19]. The situation, however, is often even more complex: neither the wave speed nor the spatiotemporal earthquake source locations are known, although sometimes a rather coarse wave speed model is available. These unknown degrees of freedom result in a mixed-integer programming problem which is harder than LSAP or QAP and challenging to solve even with a small number of earthquakes and stations [20, 21, 22].
Although the joint localization–wave speed recovery problem has a unique solution under certain conditions [7], there are no reliable algorithms that solve it provably and efficiently. Notwithstanding, there exist algorithmic attempts to perform joint full waveform (or sometimes travel time) inversion and wave speed model refinement but with temporally well-separated sources, thus implicitly avoiding the association problem (while still facing the ill-posedness) [23, 24, 25]. In the context of micro-seismicity, simple parametric models [26] and Bayesian priors [27] were used for regularization. We mention that our strategy to convert combinatorial data into measures in order to avoid combinatorial optimization is reminiscent of earlier work on unassigned Euclidean distance data [28, 29] with applications in genomics and room acoustics.
Over the last decade there has been an increasing interest in the application of machine learning in seismology, for example for phase picking [1, 2, 3, 30]. A multitude of studies addressing the phase association problem have employed either probabilistic  [31, 5, 32] or graph models [33, 34]. Some adopt a static or 1-D wave speed model, while others incorporate wave speed information indirectly through supervised learning. Since these approaches do not explicitly leverage the wave propagation physics they demand large quantities of data for training. All phenomena in the association problem are governed by the wave equation which is known, but is typically only leveraged for downstream analysis once the phases have been associated. We show that rethinking the problem in measure-theoretic terms allows one to leverage the wave propagation model from the outset.
2 Materials and Methods
2.1 Problem formulation
@@
2.2 A mixed-integer optimization problem
@@
2.3 Generative modeling of low-dimensional wave speed families
To build Harpa we assume that the wave speed models and consequently travel times exhibit a latent low-dimensional structure. We thus model the wave speeds by an autoencoder-based generative model and traveltimes by a matching continuous coordinate-based neural field [8, 9, 10]. The assumption of latent low-dimensionality is consistent with geodynamical models which generate the distribution of the elastic material properties in the interior; this has been explored for Earth’s mantle before [35, 36]. Geological random fields have been used to in conjuction with accurate earthquake simulations to study the 2019 Le Teil earthquake in a region where little geophysical data is available [37]. In fact, the wave speeds generated by our latent model ($\mathbb{R}^{L}\ni z\mapsto c(z)$) are considerably more complex than the fixed models used for association in the recent literature which only vary along the depth coordinate [38, 5, 32]. We thus define the associated travel time function as $T(s,z;r)=T_{c}(s;r)$. This enables us to rewrite the optimization problem involving $c$ as one involving the latent code $z$,
@@ 2.3 Generative modeling of low-dimensional wave speed families
2.4 Association without association
@@
In the case where some phase picks are missing or spurious picks are present, we define a loss for matching unbalanced arrivals based on sparsity-constrained optimal transport; a detailed description is given in Section 2 and results in Fig. 9c. In the 1-dimensional discrete case this loss coincides with the Wasserstein loss of the unbalanced linear assignment.
@@
where $\mathcal{E}^{*},z^{*}$ on the left hand side are the solution of (10) while $c^{*},\alpha^{*},(s^{*},\tau^{*})$ on the right hand side are the solution of (2). After obtaining $\mathcal{E}^{*}=(s^{*},\tau^{*})$ from (10), we can find $\alpha^{*}$ by solving LSAP directly, or by using only $s^{*}$ without $\tau^{*}$ by solving the TDoA problem as in (4)).
We assume that earthquakes are well-approximated by point sources which are distributed volumetrically so that we obtain sufficiently rich data. These conditions are satisfied in reality with events occurring close to ensembles of faults (or fractures) at varying orientations, or in “clouds” [39, 40].
@@
2.5 Generative travel time neural fields
@@
The way we resolve this conundrum in Harpa is as follows: We first use a deep autoencoder to build a low-dimensional latent representation of a set of complex wave speeds; see Section 3 and Fig. 3. Next, we use the fast marching method as implemented in scikit-fmm to compute the travel times between a fixed set of receivers and sources at each possible mesh node. The final and key step is then to fit a neural field (also known as an implicit neural representation) to the travel time data [9, 8]. We use a neural network $T_{\theta}$ with weights $\theta$ and periodic activation functions [8] which ensures that travel times and their derivatives with respect to source locations and wave speed latent codes can be efficiently evaluated at continuous coordinates. We train $T_{\theta}$ so that $T_{\theta^{*}}(s,z;r)\approx T(s,z;r)$ fits the travel time at a discrete set of grid points in space,
@@
The Wasserstein loss in (10) is differentiable with respect to both $z$ and $\mathcal{E}$ but it is not convex in these parameters which means that local gradient-based optimization may converge to local minima. The total number of parameters, $(3+1)\times N+L$, is small compared with common deep neural networks so there is no reason to expect that these local minima are benign. Indeed, as shown in Section 2, we observe empirically that stochastic gradient descent (SGD) often converges to poor local minima in this setting. Monte Carlo as an alternative would be prohibitively slow.
There exist modern iterative optimizers which combine the favorable aspects of sampling methods and SGD. A notable example is stochastic gradient Langevin dynamics (SGLD) [11], which can be viewed either as a mini-batch variant of Langevin dynamics or as a noisy SGD. The global convergence of Langevin dynamics can be leveraged to locate the globally-optimal earthquake and wave speed parameters which approximate their corresponding ground truth values, while benefiting from fast convergence rates [41, 42]; here the (relatively) small number of parameters helps. A SGLD iterate is computed as
@@
3 Experiments results
We now showcase the performance of Harpa on a suite of computer experiments; we work with both known and unknown wave speeds. In Section 3.2 we first test the limits of Harpa on a challenging set of unknown wave speeds in terms of the numbers of sources and receivers and show successful association in a variety of settings which were heretofore inaccessible for association. In Section 3.3 we apply Harpa to realistic wave speed models derived from the SEG / EAGE 3-D overthrust model [43] which come from a different distribution than the one used to train the latent wave speed model and show that the proposed method is robust to such perturbations. Finally, in Section 3.4 we demonstrate robustness to spurious and missing phase picks.
3.1 Confusion factor and association evaluation
@@
calculates Kendall’s $\tau$ rank correlation coefficient [44]. Succinctly, when the order of the arrivals at a station matches the order of earthquake occurrences, $\mathcal{T}$ yields a value of $1$ and $\mathsf{CF}=0$. This typically occurs during low-rate earthquakes, where the travel time is shorter than the time between the different events. Conversely, when all earthquakes occur almost simultaneously, the order of earthquake occurrence bears no correlation with the order of the phases/picks arriving at the stations. This reduces $\mathcal{T}$ closer to 00, and pushes $\mathsf{CF}$ closer to $1$, implying a higher level of confusion (the $\mathsf{CF}$ is always between 0 and 1). When studying regions such as Ridgecrest around the time of the 2019 earthquake, the magnitude cutoff is usually chosen so that the resulting CF is below 0.1 [45, 5, 32]. Fig. 4 shows representative arrival configurations for different values of $\mathsf{CF}$; Fig. 9 shows results of an exhaustive test of the proposed method at different $\mathsf{CF}$.
The difficulty of association is not solely determined by $\mathsf{CF}$; it also depends on the number of stations $N$ and the number of sources $M$ being simultaneously associated. Many algorithms initially partition the data into shorter temporal windows, usually between 5 and 20 seconds in length, with approximately 5 to 20 events in each window [3, 5]. In this proof of concept work we focus on a single window.
@@
3.2 Association via joint recovery